### v2

# Lean4Lean: Towards a Verified Typechecker for Lean, in Lean

Abstract.
In this paper we present a new “external checker” for the Lean theorem prover, written in Lean itself. This is the first complete typechecker for Lean 4 other than the reference implementation in C++ used by Lean itself, and our new checker is competitive with the original, running between 20% and 50% slower and usable to verify all of Lean’s mathlib library, forming an additional step in Lean’s aim to self-host the full elaborator and compiler. Moreover, because the checker is written in a language which admits formal verification, it is possible to state and prove properties about the kernel itself, and we report on some initial steps taken in this direction to formalize the Lean type theory abstractly and express the relation between the kernel functions and the type theory. We plan to use this project to help justify any future changes to the kernel and type theory and ensure unsoundness does not sneak in through either the abstract theory or implementation bugs.
Lean, proof assistant, external typechecker, implementation, metatheory, type theory, proof theory
ccs: Mathematics of computing Mathematical software performanceccs: Software and its engineering Formal methodsccs: Social and professional topics Systems analysis and design
1. Introduction
Lean (de Moura et al., 2015) is a theorem prover based on the Calculus of Inductive Constructions (CIC) (Paulin-Mohring, 2015), quite similar to its older brother Coq (Team, 2020), but as the name suggests, one of the ways it sought to differentiate itself was the desire to make the kernel “leaner,” relying on simpler primitives while retaining most of the power of the system. This was in particular informed by periodic news of unsoundnesses in Coq [^1] https://github.com/coq/coq/blob/master/dev/doc/critical-bugs.md due to complications of the many interacting features in the kernel. So from the beginning, Lean advertised its simpler foundations and multiple external checkers (Selsam, 2017; Ebner, 2022) as reasons that it would not suffer from the same troubles. These hopes were broadly upheld: for the entire release history of Lean 3, there were no soundness bugs reported against the kernel (written in C++), and Carneiro (Carneiro, 2019) showed the consistency of Lean with respect to ZFC with $n$ inaccessible cardinals for all $n<\omega$, so the soundness story seemed fairly strong.
But times move on, and Lean 4 now exists as an (almost) ground-up rewrite of Lean with most components now written in Lean itself. The kernel was one of the few components that was not rewritten, but it was extended with various features for performance reasons like bignum arithmetic, nested inductive types, and primitive projections, and unfortunately some soundness bugs crept in during the process and Lean’s record is no longer spotless. Another thing that was lost during the port was the external checkers: the whole metaprogramming infrastructure was redesigned so old external checkers are no longer applicable to Lean 4. While most of the new features can be treated as mere abbreviations, nested inductive types and $\eta$ for structures significantly impact the theory, and as a result the soundness proof from (Carneiro, 2019) is no longer directly applicable. (But see (Ullrich, 2023), which covers some of these modifications.)
We interpret this as a cautionary tale: It is not sufficient to see a prior system and hope to do better through sheer force of will. We are all human, and mistakes in both the theory and the implementation happen as a matter of course. The way to do better is not to rewrite, but to set up a process that structurally ensures that mistakes are either prevented or corrected. For a proof assistant, there are three main ways to do this: (1) testing, (2) independent reimplementation, and (3) formal verification. Lean has a fairly robust test suite, but testing helps more in the high-frequency low-impact domain; soundness bugs are rare and generally occur in areas that are insufficiently tested because the developer did not even consider the error condition.
The main contribution of this paper lies in (2): an independent reimplementation. To restore the soundness story of Lean to a satisfactory state, we believe it is necessary to bring back the external checkers. And what better language to do so than Lean itself? There are reasons not to use Lean for a truly independent verifier (and more external checkers are coming, in Swift [^2] https://github.com/gebner/trepplein and Rust [^3] https://github.com/ammkrn/nanoda_lib), but there are also advantages to having a kernel written in Lean:
  • The Lean elaborator is already written in Lean as a form of “whitebox automation”: users are able to make use of elaborator APIs to write their own tactics with just as much flexibility as the core language itself. Similarly, having a “whitebox kernel” makes it easier for Lean users to query the internal state and understand how Lean reduces or typechecks terms.
  • Unlike Rust and Swift, Lean is designed for both programming and proving, which means that a kernel written in Lean can have properties proved about it. And we have good reasons for wanting the kernel to satisfy certain properties…
In this paper, we will present Lean4Lean, an external verifier for Lean 4, written in Lean. The verifier itself is more than a prototype: it is fully capable of checking the entirety of Lean (the Lean compiler package), Batteries (the extended standard library), and Mathlib (the Lean mathematics library). It is not as fast as the original kernel written in C++ because the Lean compiler still has some ways to go to compete with C++ compilers, but the overhead is still reasonable in exchange for the additional guarantees it can provide, and it is suitable as a complementary step in Lean projects that wish to validate correctness in a variety of ways.
The second part of this project, which is much larger and on which we can only report partial progress, is the specification of Lean’s metatheory and verification of the Lean4Lean kernel with respect to that theory, covering mistake prevention measure (3). This is most similar to the MetaCoq project (Sozeau et al., 2020, 2019), and amounts to a formalization of the results of (Carneiro, 2019). Our principal contributions in this area are a definition of the typing judgment, along with some regularity theorems we can prove and others we can only state, as well as some invariants used in the kernel. The remainder is left as future work.
In this paper, we will avoid making value judgments about Lean or Coq and their respective theories. These systems came to be what they are now as a result of complex historical factors, and we seek only to validate the behavior of Lean as it exists, even though it lacks some type theoretic properties that would have made the job easier if true. One possible outcome of this project is a recommendation for how to improve the typing rules, but this comes at a high cost to existing users so we do not take it lightly and we have no such recommendations currently.
The paper is organized as follows: Section 2 defines the main typing judgment and describes some of its properties. Section 3 gives the core data structures of the typechecker and how they relate to the typing rules from section 2. Section 4 explains how the global structure of the environment is put together from individual typing judgments. Section 5 explains the complexities associated with supporting inductive types, treated as an extension of the base MLTT theory. Section 6 gives some performance results regarding the complete verifier, and section 7 compares this project to MetaCoq, a Coq formalization of Coq metatheory and a verified kernel.
2. Base theory
2.1. Expressions
At the heart of the theory is the VExpr [^4] The prefix “V” for “verified” disambiguates it from the Expr type used by the kernel itself. This is the more abstracted version of expressions used for specification purposes. type, which represents expressions of type theory. This is a minimalistic version of the type actually used by the kernel in section 3.1, to ease the burden of proving theoretical properties. [^5] MetaCoq seems not to do this; it uses the same type for proving metatheory and for writing metaprograms (it’s not the same type as the one Coq itself uses because that type is in ML, but it is roughly equivalent).
inductive VExpr where
| bvar (deBruijnIndex : Nat)
| sort (u : VLevel)
| const (declName : Name) (us : List VLevel)
| app (fn arg : VExpr)
| lam (binderType body : VExpr)
| forallE (binderType body : VExpr)
  • bvar \$n\$ represents the $n$’th variable in the context, counting from the inside out. As we will see, Lean itself uses “locally nameless” representation, with a combination of de Bruijn variables and named free variables, but here we only have pure de Bruijn variables.
  • sort \$u\$ represents a universe, written as Sort \$u\$ (or Prop and Type u which are syntax for Sort 0 and Sort (u+1) respectively).
  • const \$c\$ \$us\$ represents a reference to global constant $c$ in the environment, instantiated with universes $us$. (Constants in Lean can be universe-polymorphic, and they are instantiated with concrete universes at each use site.) Inductive constructors, recursors, and definitions are all represented using this constructor.
  • app \$f\$ \$a\$ is function application, written $f\;a$ or $f(a)$.
  • lam \$A\$ \$e\$ is a lambda-abstraction $\lambda x:A.\,e$. Because we are using de Bruijn variables, $x$ is not represented; instead $e$ is typechecked in an extended context where variable 00 is type $A$ and the other variables are shifted up by 1.
  • forallE \$A\$ \$B\$  [^6] The “E” suffix is used because forall is a keyword; we could use forall anyway but it would require escaping in several places. is a dependent $\Pi$-type $\Pi x:A.\,B$, also written as $\forall x:A.\,B$ because for Prop this is the same as the “for all” quantifier. It uses the same binding structure as $\lambda$.
The sort \$u\$ and const \$c\$ \$us\$ constructors depend on another type VLevel, which is the grammar of level expressions:
inductive VLevel where
| zero : VLevel
| succ : VLevel VLevel
| max : VLevel VLevel VLevel
| imax : VLevel VLevel VLevel
imax a b means (if b = 0 then 0 else max a b)
| param : Nat VLevel
These come up when typechecking expressions, for example the type of sort u is sort (succ u). The one notable case here is param \$i\$, which represents the $i$th universe variable, where $i<n$ where $n$ is the number of universe parameters to the current declaration.
This definition corresponds to the following BNF style grammar from (Carneiro, 2019), with the main change being that it is more precise about how variables and binding are represented. (For presentational reasons, we will use named variables in the informal version, but the formalization uses lift and subst functions when moving expressions between contexts.)
$\ell$ $::=0\mid\operatorname{S}\ell\mid\max(\ell,\ell)\mid\operatorname{imax}(\ell,\ell)\mid u$
$e$ $::=x\mid\mathsf{U}_{\ell}\mid c_{\bar{\ell}}\mid e\;e\mid\lambda x:e.\;e\mid\forall x:e.\;e$
$\Gamma$ $::=\cdot\mid\Gamma,x:e$
2.2. Typing and definitional equality
{mathparpagebreakable}
[…]
Figure 1. The rules for the judgment $\Gamma\vdash e\equiv e^{\prime}:\alpha$, with parameters $E$ (the global environment) and $n$ (the number of universe parameters in context), suppressed in the notation. This is VEnv.IsDefEq in the formalization.
For the typing rules, we use a slight variation of the typing judgments from (Carneiro, 2019), see Figure 1. Some remarks:
  • In (Carneiro, 2019), there were two judgments $\Gamma\vdash e:\alpha$ and $\Gamma\vdash e_{1}\equiv e_{2}$ which are mutually inductive (due to t-conv and t-proof-irrel), and $\Gamma\vdash e_{1}\equiv e_{2}:\alpha$ was a defined notion for $\Gamma\vdash e_{1}\equiv e_{2}\ \wedge\ \Gamma\vdash e_{1}:\alpha\ \wedge\ \Gamma\vdash e_{2}:\alpha$. In this version, we have only one all-in-one relation $\Gamma\vdash e_{1}\equiv e_{2}:\alpha$, and we define $\Gamma\vdash e:\alpha$ to mean $\Gamma\vdash e\equiv e:\alpha$ and define $\Gamma\vdash e_{1}\equiv e_{2}$ as $\exists\alpha.\;(\Gamma\vdash e_{1}\equiv e_{2}:\alpha)$.
    We conjecture the two formulations to be equivalent, but this version seems to be easier to prove basic structural properties about (see subsection 2.3). Moreover Lean does not have good support for mutual inductive predicates, so keeping it single-inductive makes induction proofs easier.
  • We also define $\Gamma\vdash\alpha\ \mathsf{type}$ to mean .$\exists\ell.\ (\Gamma\vdash\alpha:\mathsf{U}_{\ell})$.
  • The rules t-sort, t-const, t-extra make use of a judgment $n\vdash\ell\ \mathsf{ok}$, which simply asserts that every param \$i\$ in $\ell$ satisfies $i<n$. The $\ell\equiv\ell^{\prime}$ predicate asserts that $\ell$ and $\ell^{\prime}$ are extensionally equivalent, i.e. for all assignments $v:\mathbb{N}\to\mathbb{N}$ of natural number values to the universe variables, $\llbracket\ell\rrbracket_{v}=\llbracket\ell^{\prime}\rrbracket_{v}$.
  • The t-extra rule says that we can add arbitrary definitional equalities from the environment. This is how we will add the two supported “extensions” to the theory, for inductive types and quotients, without the details of these extensions complicating reasoning about the core theory. But we can’t actually support arbitrary definitional equalities—instead each theorem about the typing judgment has its own assumptions about what extensions are allowed. (However, note that the specific constraints on t-extra are still unfinished, pending future work on inductive type specification.)
2.3. Properties of the typing judgment
Lemma 2.1 (basic properties).
  1. (1)
    (IsDefEq.closedN) If $\Gamma\vdash e\equiv e^{\prime}:\alpha$ and $\Gamma$ is closed, then $e$, $e^{\prime}$, and $\alpha$ are closed (all free variables have indices less than $|\Gamma|$).
  2. (2)
    (IsDefEq.weakN, weakening) If $\Gamma,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha$, then $\Gamma,\Delta,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha$.
  3. (3)
    (IsDefEq.instL) If $\Gamma\vdash_{E,n}e\equiv e^{\prime}:\alpha$ and $\forall i.\;n^{\prime}\vdash\ell_{i}\ \mathsf{ok}$, then $\Gamma[\bar{u}\mapsto\bar{\ell}]\vdash_{E,n^{\prime}}e[\bar{u}\mapsto\bar{\ell}]\equiv e^{\prime}[\bar{u}\mapsto\bar{\ell}]:\alpha[\bar{u}\mapsto\bar{\ell}]$.
  4. (4)
    (IsDefEq.instN) If $\Gamma,x:\beta\vdash e_{1}\equiv e_{2}:\alpha$ and $\Gamma\vdash e_{0}:\beta$, then $\Gamma\vdash e_{1}[x\mapsto e_{0}]\equiv e_{2}[x\mapsto e_{0}]:\alpha[x\mapsto e_{0}]$.
Proof sketch.
  1. (1)
    This is a straightforward proof by induction on $\Gamma\vdash e\equiv e^{\prime}:\alpha$, except that one gets stuck at t-const because (assuming the environment is well-typed, see section 4) we know that $\vdash c_{\bar{\ell}}:\alpha$ and need that $\alpha$ is closed. So in fact this is a double induction, first over the environment $E$ and then over $\Gamma\vdash e\equiv e^{\prime}:\alpha$. We also need lemmas about $e[x\mapsto e^{\prime}]$ and $e[\bar{u}\mapsto\bar{\ell}]$ preserving closedness, but these are also direct by induction on $e$.
  2. (2)
    By induction. We use Lemma 2.1.1 in the t-const case, to show that the $\alpha$ in $\Gamma,\Gamma^{\prime}\vdash c_{\bar{\ell}}\equiv c_{\bar{\ell}^{\prime}}:\alpha$ can be used in $\Gamma,\Delta,\Gamma^{\prime}\vdash c_{\bar{\ell}}\equiv c_{\bar{\ell}^{\prime}}:\alpha$ without renaming any bound variables because $\alpha$ is closed.
  3. (3)
    By induction.
  4. (4)
    By induction on the first hypothesis. This uses
    Lemma 2.1.2 when lifting $\Gamma\vdash e_{0}:\beta$ under binders.
Lemma 2.2 (IsDefEq.defeqDF_l).
If $\Gamma,x:\alpha\vdash e_{1}\equiv e_{2}:\beta$ and $\Gamma\vdash\alpha\equiv\alpha^{\prime}:\mathsf{U}_{\ell}$, then $\Gamma,x:\alpha^{\prime}\vdash e_{1}\equiv e_{2}:\beta$.
Proof.
We have $\Gamma,x:\alpha^{\prime}\vdash x:\alpha$ by t-bvar, t-conv and weakening, and $\Gamma,\_:\alpha^{\prime},x:\alpha\vdash e_{1}\equiv e_{2}:\beta$ by weakening, so $\Gamma,x:\alpha^{\prime}\vdash e_{1}[x\mapsto x]\equiv e_{2}[x\mapsto x]:\beta[x\mapsto x]$ by Lemma 2.1.4. ∎
Lemma 2.3 (HasType.forallE_inv).
If $\Gamma\vdash(\forall x:\alpha.\;\beta):\gamma$ then $\Gamma\vdash\alpha\ \mathsf{type}$ and $\Gamma,x:\alpha\vdash\beta\ \mathsf{type}$ (and therefore also $\Gamma\vdash(\forall x:\alpha.\;\beta)\ \mathsf{type}$).
Proof sketch.
By induction on $\Gamma\vdash e_{1}\equiv e_{2}:\gamma$, assuming that one of $e_{1}$ or $e_{2}$ is $\forall x:\alpha.\;\beta$. Most cases are trivial or inapplicable. Of those that remain:
  • t-all: If $e_{1}$ is $\forall x:\alpha.\;\beta$ we are done; if $e_{2}$ is $\forall x:\alpha.\;\beta$ then we obtain $\Gamma,x:\alpha^{\prime}\vdash\beta\ \mathsf{type}$ and need Lemma 2.2 to get $\Gamma,x:\alpha\vdash\beta\ \mathsf{type}$.
  • t-beta: For this to apply, it must be that we have $(\lambda y:\delta.\;e)\;e^{\prime}\equiv e[y\mapsto e^{\prime}]$ where $e[y\mapsto e^{\prime}]$ is $\forall x:\alpha.\;\beta$. So either $e=y$ and $e^{\prime}=\forall x:\alpha.\;\beta$, in which case the inductive hypothesis for $e^{\prime}$ applies, or $e=\forall x:\alpha^{\prime}.\;\beta^{\prime}$ with $\alpha^{\prime}[y\mapsto e^{\prime}]=\alpha$ and $\beta^{\prime}[y\mapsto e^{\prime}]=\beta$ in which case $\Gamma,y:\delta\vdash\alpha^{\prime}\ \mathsf{type}$ and $\Gamma,y:\delta,x:\alpha^{\prime}\vdash\beta^{\prime}\ \mathsf{type}$ by the inductive hypothesis for $e$, and Lemma 2.1.4 applies.
  • t-extra: It could be that $e[\bar{u}\mapsto\bar{\ell}]$ is $\forall x:\alpha.\;\beta$, but this can only happen if $e=\forall x:\alpha^{\prime}.\;\beta^{\prime}$ with $\alpha^{\prime}[\bar{u}\mapsto\bar{\ell}]=\alpha$ and $\beta^{\prime}[\bar{u}\mapsto\bar{\ell}]=\beta$, so by induction hypothesis (for the environment) $\vdash\alpha^{\prime}\ \mathsf{type}$ and $x:\alpha^{\prime}\vdash\beta^{\prime}\ \mathsf{type}$, and we conclude using level substitution and weakening.
Lemma 2.4 (HasType.sort_inv).
If $\Gamma\vdash_{E,n}\mathsf{U}_{\ell}:\gamma$ then $n\vdash\ell\ \mathsf{ok}$ (and therefore also $\Gamma\vdash\mathsf{U}_{\ell}:\mathsf{U}_{\operatorname{S}\ell}$).
Proof sketch.
Similar to Lemma 2.3. ∎
Theorem 2.5 (IsDefEq.isType).
If $\vdash\Gamma\ \mathsf{ok}$ and $\Gamma\vdash e:\alpha$ then $\Gamma\vdash\alpha\ \mathsf{type}$.
Proof sketch.
By induction, using weakening and substitution lemmas. The nontrivial cases are:
  • t-app: We have $\Gamma\vdash(\forall x:\alpha.\;\beta)\ \mathsf{type}$ from the IH and $\Gamma\vdash e_{2}:\alpha$ by assumption, and from Lemma 2.3 we get $\Gamma,x:\alpha\vdash\beta\ \mathsf{type}$, so $\Gamma\vdash\beta[x\mapsto e_{2}]\ \mathsf{type}$ by the substitution lemma.
  • t-all: We have $\Gamma\vdash\mathsf{U}_{\ell_{1}}\ \mathsf{type}$ and $\Gamma,x:\alpha\vdash\mathsf{U}_{\ell_{1}}\ \mathsf{type}$ from the IH, so by Lemma 2.4 we have $n\vdash\ell_{1},\ell_{2}\ \mathsf{ok}$, therefore $\Gamma\vdash\mathsf{U}_{\operatorname{imax}(\ell_{1},\ell_{2})}\ \mathsf{type}$.
Lemma 2.6 (IsDefEq.instDF).
If $\Gamma,x:\alpha\vdash f\equiv f^{\prime}:\beta$ and $\Gamma\vdash a\equiv a^{\prime}:\alpha$ then $\Gamma\vdash f[x\mapsto a]\equiv f^{\prime}[x\mapsto a^{\prime}]:\beta[x\mapsto a]$.
Proof sketch.
This is a generalization of Lemma 2.1.4, but we need Theorem 2.5 to prove that $\alpha$ and $\beta$ are well-typed to apply the rules below.
First we show the claim assuming $\Gamma\vdash\beta[x\mapsto a]\equiv\beta[x\mapsto a^{\prime}]:\mathsf{U}_{\ell}$. In this case we have
$f[x\mapsto a]$ $\equiv(\lambda x:\alpha.\;f)\;a$
$\equiv(\lambda x:\alpha.\;f^{\prime})\;a^{\prime}$
$\equiv f^{\prime}[x\mapsto a^{\prime}]\qquad{}:\beta[x\mapsto a],$
where we use t-beta twice, and use the assumption $\beta[x\mapsto a]\equiv\beta[x\mapsto a^{\prime}]$ to justify the last step ecause t-beta gives the equality at the type $\beta[x\mapsto a^{\prime}]$ instead.
To finish, we apply the lemma twice, once with $f$ and $\beta$ so that it suffices to show $\beta[x\mapsto a]\equiv\beta[x\mapsto a^{\prime}]:\mathsf{U}_{\ell}$ and then again with $\beta$ in place of $f$ and $\mathsf{U}_{\ell}$ in place of $\beta$ so that it suffices to show $\mathsf{U}_{\ell}[x\mapsto a]\equiv\mathsf{U}_{\ell}[x\mapsto a^{\prime}]:\mathsf{U}_{\operatorname{S}\ell}$, which is true by reflexivity. ∎
2.4. Conjectured properties of the typing judgment
Unfortunately, there are more structural properties of the typing judgment beyond the results in the previous section. Many of these are close relatives of each other and can be proved from other conjectures in this section.
Conjecture 2.7 (Unique typing).
If $\Gamma\vdash e:\alpha$ and $\Gamma\vdash e:\beta$, then $\Gamma\vdash\alpha\equiv\beta:\mathsf{U}_{\ell}$ for some $\ell$.
This has a major simplifying effect on the theory. Here are some consequences (with proofs omitted, but they are simple algebraic consequences of earlier lemmas):
Corollary 2.8 (Consequences of unique typing).
  1. (1)
    (IsDefEqU.trans) If $\Gamma\vdash e_{1}\equiv e_{2}$ and $\Gamma\vdash e_{2}\equiv e_{3}$, then $\Gamma\vdash e_{1}\equiv e_{3}$. (This is t-trans but without requiring the types to match.)
  2. (2)
    (isDefEq_iff) $\Gamma\vdash e\equiv e^{\prime}:\alpha$ if and only if $\Gamma\vdash e:\alpha$, $\Gamma\vdash e^{\prime}:\alpha$, and $\Gamma\vdash e\equiv e^{\prime}$. (Hence this formulation implies the one of (Carneiro, 2019).)
  3. (3)
    (IsDefEqU.defeqDF) If $\Gamma\vdash e\equiv e^{\prime}:\alpha$ and $\Gamma\vdash\alpha\equiv\beta$, then $\Gamma\vdash e\equiv e^{\prime}:\beta$. (This is t-conv but with an untyped definitional equality.)
This group of theorems is proved mutually with Conjecture 2.7 in (Carneiro, 2019):
Conjecture 2.9 (Definitional inversion).
  1. (1)
    (IsDefEqU.sort_inv) If $\Gamma\vdash\mathsf{U}_{\ell}\equiv\mathsf{U}_{\ell^{\prime}}$ then $\ell\equiv\ell^{\prime}$.
  2. (2)
    (IsDefEqU.forallE_inv) If $\Gamma\vdash(\forall x:\alpha.\;\beta)\equiv(\forall x:\alpha^{\prime}.\;\beta^{\prime})$ then $\Gamma\vdash\alpha\equiv\alpha^{\prime}$ and $\Gamma,x:\alpha\vdash\beta\equiv\beta^{\prime}$. [^7] This theorem is sometimes called “injectivity of $\forall$” but this name is slightly misleading because $\forall$ is not injective as a function on types, which is to say if we use propositional equality instead of def.eq then the theorem fails.
  3. (3)
    (IsDefEqU.sort_forallE_inv) $\Gamma\vdash\mathsf{U}_{\ell}\not\equiv(\forall x:\alpha.\;\beta)$.
The reason Conjecture 2.7 and Conjecture 2.9 have been downgraded from theorems in (Carneiro, 2019) to conjectures here is because the proof has an error in one of the technical lemmas, and it remains to be seen if it is possible to salvage the proof. More precisely, the proof constructs a stratification $\vdash_{i}$ of the typing judgment in order to break the mutual induction between typing and definitional equality, but this stratification does not and cannot respect substitution; that is, if $\Gamma,x:\beta\vdash_{i}e:\alpha$ and $\Gamma\vdash_{j}e^{\prime}:\beta$, then the proof requires $\Gamma\vdash_{\max(i,j)}e[x\mapsto e^{\prime}]:\alpha$ but only $\Gamma\vdash_{i+j}e[x\mapsto e^{\prime}]:\alpha$ holds.
We still have reasons to believe the conjectures here are true, but more work is needed to determine how to structure the proof by induction. The proof of soundness is not impacted because there are alternative routes to construct the model that avoid unique typing (also described in (Carneiro, 2019)), but these conjectures are necessary in at least some form in order to prove the correctness of the typechecker (see section 3).
Another class of conjectured theorems concerns the “invertibility” of weakening:
Conjecture 2.10 (Strengthening).
If $e,e^{\prime}$ do not mention any variables in $\Delta$, then $\Gamma,\Delta,\Gamma^{\prime}\vdash e\equiv e^{\prime}$ if and only if $\Gamma,\Gamma^{\prime}\vdash e\equiv e^{\prime}$.
Corollary 2.11 (Consequences of strengthening).
  1. (1)
    (IsDefEq.skips) If $\Gamma,\Delta,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha$ and $e,e^{\prime}$ do not mention the variables in $\Delta$, then there exists $\alpha^{\prime}$ not mentioning $\Delta$ such that $\Gamma,\Delta,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha^{\prime}$.
  2. (2)
    (IsDefEq.weakN_iff) If $e,e^{\prime},\alpha$ do not mention any variables in $\Delta$, then $\Gamma,\Delta,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha$ if and only if $\Gamma,\Gamma^{\prime}\vdash e\equiv e^{\prime}:\alpha$.
  3. (3)
    (OnCtx.weakN_inv) If $\vdash\Gamma,\Delta,\Gamma^{\prime}\ \mathsf{ok}$ and $\Gamma^{\prime}$ does not mention the variables in $\Delta$ then $\vdash\Gamma,\Gamma^{\prime}\ \mathsf{ok}$.
Note that in the formalization, the “$e$ does not mention the variables in $\Delta$” clauses are expressed by saying that $e$ is the lift of some $e^{\prime}$ in the smaller context, because of the use of de Bruijn variables.
3. The typechecker
3.1. Relating VExpr to Expr
The actual Lean kernel does not use any of the machinery from the previous section directly. Instead, it works with another type, Expr:
inductive Expr where
| bvar (deBruijnIndex : Nat)
| fvar (fvarId : FVarId)
| mvar (mvarId : MVarId)
| sort (u : Level)
| const (declName : Name) (us : List Level)
| app (fn arg : Expr)
| lam (binderName : Name) (type body : Expr)
(info : BinderInfo)
| forallE (binderName : Name) (type body : Expr)
(info : BinderInfo)
| letE (declName : Name) (type value body : Expr)
(nonDep : Bool)
| lit : Literal Expr
| mdata (data : MData) (expr : Expr)
| proj (typeName : Name) (idx : Nat)
(struct : Expr)
Comparing this with VExpr reveals some differences:
  • mvar \$n\$ is for metavariables, which are used for incomplete proofs in the elaborator. It does not matter much for us because the kernel does not create them or allow them in terms submitted to it.
  • fvar \$n\$ is a “free variable”, which is identified by a name rather than as an index which changes depending on the context. Although these are also not allowed in terms provided to the kernel, it does create fvar \$n\$ expressions during typechecking, and converting between locally nameless and pure de Bruijn is one of the major differences we will have to deal with in verification of the kernel.
  • The Level type (not shown) used in the sort constructor also differs in similar ways to VLevel, containing an mvar constructor we don’t care about and a param constructor which takes a Name instead of a Nat.
  • The lam and forallE constructors contain information which is relevant only for elaboration and printing like the names of bound variables and whether the argument is implicit or explicit.
  • The letE construct is used for let x := e1; e2 expressions. We handle these by simply expanding them to $e_{2}[x\mapsto e_{1}]$ in the abstract theory, meaning that we don’t have to handle this constructor while doing induction proofs about the typing judgment. The kernel puts these in the context and unfolds them as necessary.
  • The lit constructor is for literals:
    inductive Literal where
    | natVal (val : Nat)
    | strVal (val : String)
    This is one of the new features of the Lean 4 kernel: natural numbers are represented directly as bignums and several functions like Nat.mul are overridden with a native implementation rather than using the inductive structure of Nat directly (effectively writing numerals in unary). For the purpose of modeling in VExpr however, we just perform exactly this (exponential-size) unfolding in terms of Nat.zero and Nat.succ applications. Similarly, String is a wrapper around List Char for the purpose of modeling, even though the actual representation in the runtime is as a UTF-8 encoded byte string.
  • The mdata \$d\$ \$e\$ constructor is equivalent to $e$ for modeling purposes (this is used for associating metadata to expressions). We can’t completely ignore it though because it has a different hash than $e$, which impacts some of the caches used in the kernel.
  • proj \$c\$ \$i\$ \$e\$ is a projection out of a structure or
    structure-like inductive type. (See section 5.) This is the most challenging constructor to desugar, because even though it is equivalent to an application of the recursor for the inductive type, we need the type of $e$ to construct this term. This means that the desugaring function must be type-aware, and is only well-defined up to definitional equality.
It is worth mentioning that Expr was not defined as part of this project — this is a type imported directly from the Lean elaborator. In fact, the C++ Lean kernel uses FFI (foreign-function interface) to Lean in order to make use of this type, so Lean4Lean and the C++ kernel are sharing this data structure.
As a result of these considerations, we use an inductive type to encapsulate the translation from Expr to VExpr:
inductive TrExpr (env : VEnv) (Us : List Name) :
VLCtx Expr VExpr Prop
[…]
The type VLCtx appearing here is new, and it contains the information we need to translate between the Expr and VExpr types:
inductive VLocalDecl where
| vlam (type : VExpr) x : type
| vlet (type value : VExpr) x : type := value
def VLCtx := List (Option FVarId × VLocalDecl)
[…]
def toCtx : VLCtx List VExpr
| (_, vlam α) :: Δ => α :: toCtx Δ
| (_, vlet _ _) :: Δ => toCtx Δ
We have to be careful to reindex the Expr.bvar \$i\$ indices though, since this includes both let- and lambda-variables while VExpr.bvar \$i\$ only counts the lambda-variables, with the let-variables expanded to terms.
The Option FVarId in the context represents the “name” of the variables. Lean follows the “locally nameless” discipline, in which most operations act only on “closed terms” (where closed means that there are no bvar \$i\$ variables outside a binder), so any time it needs to process a term under a binder it first instantiates the term, replacing bvar \$(i+1)\$ with bvar \$i\$ and bvar 0 with fvar \$a\$, where \$a\$ : FVarId is a (globally) fresh variable name. We store these names in $\Delta$ so that we know how to relate them to de Bruijn indices in the context. The name generator uses a counter in the global state of the typechecker, and there is an invariant that all FVarIds in the state are less than the counter value.
The reason it is an Option is because when we are typing an open term, we still have to deal with variables that have not been assigned FVarIds. Even though Lean does not directly handle such terms, they appear as subterms of expressions that are handled so we need both cases to have a compositional specification.
3.2. The typechecker implementation
The kernel itself is defined completely separately from VExpr, and is a close mirror of the C++ code, modulo translation into functional style. For example, the code for inferring/checking the type of $\lambda x:\alpha.\;e$:
Lean4Lean
def inferLambda (e : Expr) (inferOnly : Bool) :
RecM Expr := loop #[] e where
loop fvars : Expr RecM Expr
| .lam name dom body bi => do
let d := dom.instantiateRev fvars
let id := ⟨← mkFreshId
let lc := (← getLCtx).mkLocalDecl id name d bi
withLCtx lc do
let fvars := fvars.push (.fvar id)
if !inferOnly then
_ ensureSortCore
(← inferType d inferOnly) d
loop fvars body
| e => do
let r inferType
(e.instantiateRev fvars) inferOnly
let r := r.cheapBetaReduce
return (← getLCtx).mkForall fvars r
// C++ Lean kernel
expr type_checker::infer_lambda(expr const & _e,
bool infer_only) {
flet<local_ctx> save_lctx(m_lctx, m_lctx);
buffer<expr> fvars;
expr e = _e;
while (is_lambda(e)) {
expr d = instantiate_rev(binding_domain(e),
fvars.size(), fvars.data());
expr fvar = m_lctx.mk_local_decl(m_st->m_ngen,
binding_name(e), d, binding_info(e));
fvars.push_back(fvar);
if (!infer_only) {
ensure_sort_core(
infer_type_core(d, infer_only), d); }
e = binding_body(e); }
expr r = infer_type_core(instantiate_rev(
e, fvars.size(), fvars.data()), infer_only);
r = cheap_beta_reduce(r);
return m_lctx.mk_pi(fvars, r); }
Although control flow is expressed differently (here, the replacement of a while loop with a tail-recursive function), the two implementations do essentially the same things in the same order. As a result, most bugs or peculiarities we have discovered in the course of formalizing Lean4Lean are replicated in the C++ kernel, which gives us hope that verifying Lean4Lean will lend strong evidence to the correctness of the C++ code as well. We have taken care to write things in a style which will lend itself well to the verification task.
3.2.1. Termination and the RecM monad
The function above is defined in the RecM monad, which is defined like so:
structure Methods where
isDefEqCore : Expr Expr M Bool
whnfCore (e : Expr) (cheapRec := false)
(cheapProj := false) : M Expr
whnf (e : Expr) : M Expr
inferType (e : Expr) (inferOnly : Bool) : M Expr
abbrev RecM := ReaderT Methods M
This is a monad stack on top of the main TypeChecker.M monad which contains the actual state of the typechecker. The use of Methods here is a trick for cutting the knot of mutual definitions involved in the kernel: removing calls to these four functions makes the call graph acyclic (except for self-loops corresponding to directly recursive definitions). So we can avoid mutual definitions — every definition only refers to itself and earlier functions.
There are a few reasons to avoid mutual definitions:
  • Lean is bad at elaborating large mutual blocks, because it cannot process them sequentially. This is not critical but it makes working on the code more difficult.
  • Mutual definitions generally compile to a well-founded recursion, while we want to utilize structural recursion when possible because of the better induction principles.
  • Theorems about mutual definitions also have to be mutual, so this has a tendency toward lack of modularity.
  • As we will see, there is no suitable well-founded measure in the first place, so we would be forced to declare the definition as partial, but we absolutely cannot do this if we want any hope of proving correctness because it would make our functions opaque, so we need another option.
Having a specific list of cut points for the graph of mutual definitions doesn’t make them any less recursive, but at least gives us a clear structure in which to perform the proof. The four functions above were found by computer search, but they are sensible functions with a clear spec so it is not too surprising that they are used throughout the kernel:
  • isDefEqCore \$s\$ \$t\$ returns true if $\Gamma\vdash s\equiv t$ is provable. (This is normally used via isDefEq \$s\$ \$t\$, which also caches this result, but some callers use isDefEqCore directly.)
  • whnf \$e\$ returns the weak head normal form (WHNF) of $e$. From a modeling perspective, the main important property is that if it returns $e^{\prime}$ and $e$ is well-typed then $\Gamma\vdash e\equiv e^{\prime}$ is provable.
  • whnfCore \$e\$ cheapRec cheapProj has the same specification as whnf \$e\$, but it lacks some of the early exit paths. cheapRec and cheapProj are flags affecting which kinds of terms are unfolded. The reason both whnf and whnfCore show up separately in this list is because they are both used in other functions (i.e. some functions call whnfCore directly because e.g. the fast path isn’t applicable).
  • inferType \$e\$ inferOnly infers or typechecks a term. That is, if inferType \$e\$ inferOnly returns $\alpha$ then $\Gamma\vdash e:\alpha$ holds if either $e$ is well typed or inferOnly is false.
Now, this method of calling functions to avoid recursion only kicks the problem up one step. How do we construct an element of Methods if all the functions require another Methods? Ideally, we would actually prove the termination of the kernel, because DTT is supposed to be terminating. However:
  • It is unlikely that we can prove termination of a typechecker for Lean in Lean, because although our soundness proof does not depend on termination, MetaCoq’s does (Sozeau et al., 2019), and generally termination measures for DTT require large cardinals of comparable strength to the proof theory. We are up against Gödel’s incompleteness theorem, so anything that would imply the unconditional soundness of Lean won’t be directly provable.
  • Besides this, the Lean type theory is known not to terminate. Coquand and Abel (Abel and Coquand, 2020) constructed a counterexample to strong normalization using reduction of proofs, and this can be shown to impact definitional equality checks even for regular types:
    /-! Andreas-Abel construction of nontermination in proofs -/
    def True := p : Prop, p p
    def om : True := fun A a =>
    @cast (True True’) A
    (propext fun _ => a, fun _ => id⟩)
    (fun z => z (True True’) id z)
    def Om : True := om (True True’) id om
    #reduce Om whnf nontermination
    /-! nontermination outside proofs: -/
    inductive Foo : Prop | mk : True Foo
    def foo : Foo := Om _ (Foo.mk fun _ => id)
    example : foo.recOn (fun _ => 1) = 1 := by
    rfl isDefEq nontermination
    Essentially, this is a combination of impredicativity, proof irrelevance, and subsingleton elimination.
  • The kernel does not loop forever in many cases because this is a bad user experience — it has timeouts and depth limits. Not all parts of the kernel have such limits, but it does give us a reasonable design principle which fortuitously solves our termination problem. [^8] There are alternative methods for representing partial functions such as the Bove-Capretta method (take a proof of termination as argument), but we do not want to use this as the kernel API is fixed to match Lean’s actual kernel function, and we also actually want it to time out on extreme cases.
So we use what is arguably the standard solution for defining partial functions in a language like Lean or Coq: use a fuel parameter, a natural number which counts the number of nested recursive calls to one of the Methods, and throw a deepRecursion error if we run out of fuel. Currently, this limit is a fixed constant (1000), which turns out to be sufficient for checking all of Mathlib, but this could be made configurable. (Note that this is not a limit on the depth of expressions exactly, these use structural recursion and hence need no fuel for the termination argument; instead fuel is only consumed when making a recursive call that is not otherwise decreasing, for example when reducing definitions to WHNF. The code that is most likely to hit depth limits is in proofs by reflection, but these are comparatively rare, in part because the kernel algorithm for this is not very efficient.)
4. The global environment
The environment is the global structure that ties together individual declarations. It has made some appearances in the previous sections already, because the environment is needed to typecheck constants as well as definitional extensions, given in the t-const and t-extra rules.
The Lean type for this, Environment, is complex and contains many details irrelevant to the kernel, but luckily we only really care about a few operations on it:
  • add : Environment ConstantInfo Environment
    – adds a ConstantInfo declaration to the environment
  • mkEmptyEnvironment : IO Environment
    – constructs an environment with no constants [^9] This is in IO because environment extensions can perform computation in IO for their initialization. But the constant map, which is what we care to verify, is initialized to the empty map.
  • find? : Environment Name Option ConstantInfo
    – retrieves a declaration by name
In effect, the environment is just a fancy hashmap which indexes ConstantInfo declarations by their names.
The corresponding theory type is called VEnv, with the definition:
structure VConstant where
(uvars : Nat) (type : VExpr)
structure VDefEq where
(uvars : Nat) (lhs rhs type : VExpr)
structure VEnv where
constants : Name Option (Option VConstant)
defeqs : VDefEq Prop
This is a very simple type, essentially exactly what is needed to satisfy the requirements of the typing judgment.
  • constants maps a name to a constant, if defined, where a constant is given by its universe arity and its type. It returns none if a constant by that name does not exist, and some none if the name has been “blocked”, meaning that there is no constant there but we still want to make it illegal to make a definition with that name. This is how we model unsafe declarations when typechecking safe declarations: they “don’t exist” for most purposes, but it is nevertheless not allowed to shadow them with another safe declaration. This corresponds to the $\bar{u}.(c_{\bar{u}}:\alpha)$ declaration appearing in t-const.
  • defeqs is a set of (anonymous) VDefEqs, containing two expressions to be made definitionally equal and their type. This corresponds to the $\bar{u}.(e\equiv e^{\prime}:\alpha)$ declaration appearing in t-extra.
In some sense this doesn’t actually answer many interesting questions about where the constants come from or what definitional equalities are permitted. For that, we need VDecl, the type of records that can be used to update the environment. The declarations are:
  • block \$n\$ blocks constant $n$, setting env.constants \$n\$ = some none and preventing a redeclaration as mentioned above.
[…]
    adds $c:\alpha$ as an axiom to the constant map.
[…]
    checks that $e:\alpha$ and then leaves the environment as is.
  • induct (\$d\$ : VInductDecl) adds all the constants and defeqs from an inductive declaration (see section 5).
  • quot adds the quotient axioms and definitional equality. This is a type operator $\alpha/R$ where $R:\alpha\to\alpha\to\mathtt{Prop}$ with $\mbox{mk}:\alpha\to\alpha/R$ and $\mbox{lift}:(f:\alpha\to\beta)\to(\forall xy.\;R\;x\;y\to f\;x=f\;y)\to(\alpha/R\to\beta)$, with the property that $(\mbox{lift}\;f\;h\;(\mbox{mk}\;a)\equiv f\;a:\beta)$. (One can construct every part of this in Lean except for the definitional equality.)
This enumeration is fairly similar to the Declaration type which is the actual front end to the kernel, but it lacks unsafe declarations and mutualDefnDecl (which, despite the name, is not the way mutual definitions are sent to the kernel, but rather represents unsafe declarations with unchecked self-referential definitions). For the most part we do not attempt to model unsafe declarations because these are typechecked with certain checks disabled, and this makes them no longer follow the theory, but this is deliberate. There isn’t much we can say about such definitions since they can violate type safety and cause undefined behavior, and in any case they don’t play a role in checking theorems.
5. Inductive types
Lean’s implementation of inductive types is based on Dybjer (Dybjer, 1994). It notably differs from Coq in that rather than having primitive fix and match expressions, we have constants called “recursors” generated from the inductive specification. This choice significantly simplifies the VExpr type, which as we have seen contains no special support for anything inductive-related except for the generic def.eq. hook t-extra. It also means we do not need a complex guard checker, something which MetaCoq currently axiomatizes (Sozeau et al., 2019) and which has been the source of soundness bugs in the past. Instead, the complexity is pushed to the generation of the recursor for the inductive type. [^10] It would be premature to draw conclusions about the relative complexity of these two approaches though, as many of the important theorems are unfinished.
For single inductives, the algorithm is described in full in (Carneiro, 2019). In Lean 3, nested and mutual inductives were simulated using single inductives so the kernel only had to deal with the single inductive case. However this simulation process was both costly and did not cover the full grammar of nested and mutual inductives, so Lean 4 moved the checking to the kernel.
For a simple example of a mutual inductive, we can define Even and Odd by mutual induction like so:
mutual
inductive Even : Nat Prop where
| zero : Even 0
| succ : Odd n Even (n+1)
inductive Odd : Nat Prop where
| succ : Even n Odd (n+1)
end
which generates the definitions:
Even : Nat Type
Even.zero : Even 0
Even.succ : {n : Nat}, Odd n Even (n + 1)
Odd : Nat Type
Odd.succ : {n : Nat}, Even n Odd (n + 1)
Even.rec.{u} :
{motive_1 : (a : Nat) Even a Sort u}
{motive_2 : (a : Nat) Odd a Sort u}
motive_1 0 Even.zero
({n : Nat} (a : Odd n) motive_2 n a motive_1 (n + 1) (Even.succ a))
({n : Nat} (a : Even n) motive_1 n a motive_2 (n + 1) (Odd.succ a))
{a : Nat} (t : Even a) motive_1 a t
Odd.rec.{u} :
same as Even.rec
{a : Nat} (t : Odd a) motive_2 a t
and the definitional equalities:
@Even.rec m1 m2 Fz Fs Fs 0 Even.zero Fz
@Even.rec m1 m2 Fz Fs Fs (n+1) (Even.succ e)
@Fs n e (@Odd.rec m1 m2 Fz Fs Fs n e)
@Odd.rec m1 m2 Fz Fs Fs (n+1) (Odd.succ e)
@Fs n e (@Even.rec m1 m2 Fz Fs Fs n e)
For nested inductives, the checking procedure effectively amounts to running the simulation procedure from Lean 3 to get a mutual inductive type and performing well-formedness checks there, then removing all references to the unfolded type in the recursor. For example:
inductive T : Type where
| mk : List T T
T.rec.{u} :
{m1 : T Sort u}
{m2 : List T Sort u}
((a : List T) m2 a m1 (T.mk a))
m2 []
((head : T) (tail : List T)
m1 head m2 tail m2 (head :: tail))
(t : T) m1 t
This recursor (there is also T.rec_1, not shown) appears exactly like one would expect for a mutual inductive with T and List_T defined the same way as List, except that all occurrences of List_T in the recursor type have been removed in favor of List T.
Currently, Lean4Lean contains a complete implementation of inductive types, including nested and mutual inductives and all the typechecking effects of this, but not much work has been done on the theoretical side, defining what an inductive specification should generate. Ideally we would like to do it in a way which is not so intimately tailored to the nested inductive collection process (i.e. we can see the order in which subterms are visited by the order of motive_\$n\$ arguments in the above recursor) and instead allows a range of implementation strategies which can all be proved sound against a set-theoretic model.
5.1. Eta for structures
Another new kernel feature in Lean 4 is primitive projections, and eta for structures, which says that if $s$ is a structure (which is to say, a one-constructor non-recursive inductive type), then $s\equiv\mbox{mk}(s.1,\dots,s.n)$ where $s.i$ is the $i$’th projection (Expr.proj). This was already true as a propositional equality proven by induction on $s$, but it is now a definitional equality and the kernel has to unfold structures when needed to support this. This does not fundamentally change the theory; indeed in (Carneiro, 2019) a mild form of eta for pairs was introduced in order to make the reduction to W-types work, so we believe that the main results should still hold even with this extension.
6. Results
Lean4Lean contains a command-line frontend, which can be used to validate already-compiled Lean projects. It operates on .olean files, which are essentially serialized (unordered) lists of ConstantInfo objects, and the frontend topologically sorts these definitions and passes them to the Lean4Lean kernel, which is a function
def addDecl (env : Environment) (decl : Declaration) (check := true) :
Except KernelException Environment
that is a drop-in replacement for Environment.addDecl, which is an opaque Lean function implemented by FFI (foreign-function interface) to the corresponding C++ kernel function.
This is the same interface (and to some extent, the same code) as lean4checker, [^11] https://github.com/leanprover/lean4checker which does the exact same thing but calls addDecl instead, effectively using the C++ kernel as an external verifier for Lean proofs. It may seem odd to use the Lean kernel as an external verifier for itself, but this can catch cases where malicious (or confused) Lean code makes use of the very powerful metaprogramming framework to simply bypass the kernel and add constants to the environment without typechecking them.
For our purposes, it acts as a very handy benchmark for comparison, since we are performing the same operations but with a different kernel. Running lake env lean4lean --fresh Mathlib in the mathlib4 project will check the Mathlib module and all of its dependencies. (This is running in single-threaded mode; it can also check modules individually, assuming the correctness of imports, but it runs into memory usage limitations so is harder to benchmark reliably.)
lean4checker lean4lean ratio
Lean 37.01 s 44.61 s 1.21
Batteries 32.49 s 45.74 s 1.40
Mathlib (+ B. + Lean) 44.54 min 58.79 min 1.32
Figure 2. Comparison of the original C++ implementation of the Lean kernel (lean4checker) with lean4lean (Lean) on major Lean packages. Tests were performed on a 12 core 12th Gen Intel i7-1255U @ 2.1 GHz, single-threaded, on rev. 526c94c of mathlib4.
The results are summarized in Figure 2. The new Lean implementation is around 30% slower than the C++ implementation, which we attribute mainly to shortcomings in the Lean compiler and data representation compared to C++. Nevertheless, it is within an order of magnitude and practically usable on large libraries, which we consider a strong sign. We do not expect this to get any faster without Lean compiler improvements, but we also intend to keep the speed from regressing as we add more verification. Because the verification is external to the executable itself, this is possible in theory, but we may need to create new verification tooling to make it doable in practice.
7. Related work & Conclusion
Self-verification of ITP systems has been done in Coq (Barras and Werner, 1997), HOL Light (Harrison, 2006), Isabelle (Nipkow and Roßkopf, 2021), and Metamath Zero (Carneiro, 2020), and to some extent self-verification also overlaps with bootstrapping theorem provers such as Milawa (Davis, 2009) and Candle (Abrahamsson et al., 2022; Kumar et al., 2014). But the most similar work is unquestionably MetaCoq (Sozeau et al., 2020, 2019, 2023), which is a project to develop a verified typechecker for Coq, in Coq. It is significantly more complete than the previous effort (Barras and Werner, 1997), but is not yet capable of verifying the Coq standard library due to lack of eta-conversion and template polymorphism; eta-conversion is work in progress (Lennon-Bertrand, 2022b), and universe polymorphism is planned to be reorganised in Coq’s kernel. With Lean4Lean we took the approach of quickly getting to feature parity with the real Lean kernel and proving correctness later, so we ended up with a feature-complete typechecker but are still lacking in some theorems. Our work also overlaps to some extent with Template-Coq (Anand et al., 2018; Boulier, 2018; Winterhalter, 2020), a certified metaprogramming framework for Coq developed in the MetaCoq repo.
As was mentioned, MetaCoq also has to deal with many of the same issues proving properties of the typing judgment, and one may hope for proofs from MetaCoq to be useful in Lean4Lean and vice-versa. Unfortunately, there are some important details that differ:
  • In MetaCoq, the conversion relation is a partial order rather than an equivalence relation, because of universe cumulativity, so Conjecture 2.7 doesn’t hold as stated. Nevertheless, there are theorems regarding “principal types” which can be used to play the same role.
  • Unlike in Lean4Lean, the conversion relation is untyped: there is no mutual induction between typing and definitional equality, which massively simplifies matters. In Lean this is unfortunately not an option because of the t-proof-irrel rule, which is absent in Coq.
  • Lean’s typechecker is known to not be complete, owing to some of the undecidability results in (Carneiro, 2019), but MetaCoq (Sozeau et al., 2023; Lennon-Bertrand, 2022a) has a proof of completeness.
In this paper we presented a new external typechecker for Lean, but it is also the beginning of a larger project to both verify the correctness of a typechecker with respect to the Lean metatheory, and also to verify that the Lean metatheory is consistent relative to well-studied axiomatic foundations like ZFC with inaccessible cardinals. Lean 4 is written in Lean, so this is a prime opportunity to have a production-grade ITP which is able to bootstrap at the logical level.
Acknowledgements.
I would like to thank Jeremy Avigad for their support and encouragement, Yannick Forster and the anonymous reviewers for reading early drafts of this work, and Matthieu Sozeau and the rest of the MetaCoq team for collaborating and taking interest in this work. Work supported by the Hoskinson Center for Formal Mathematics at Carnegie Mellon.